OpenStudy (anonymous):

1. In the function, A(t)=4(1/2)^1/5t A(t) represents the: initial amount of the substance growth factor number of times to complete a cycle amount of the substance present at time t 2. In the function, M(t)=8(2)^1/6t 6 represents the: initial amount of the substance growth factor number of times to complete a cycle amount of the substance present at time t

OpenStudy (anonymous):

Did that help? it appears there is two of you working on this

OpenStudy (anonymous):

Not really :(

OpenStudy (anonymous):

I think I'm a bit more behind than she is

OpenStudy (anonymous):

at an initial amount what is the time?

OpenStudy (anonymous):

amount of the substance present at time t ?

OpenStudy (anonymous):

alright i'll use the same analogy. If you had 3 gummy bears at first before doing anything (meaning the time was 0) and then you gave 2 gummy bears an hour later (t=1), what is another name for when you had 3 gummy bears

OpenStudy (anonymous):

Initial time

OpenStudy (anonymous):

yes that would be the initial time but what would the gummy bears be

OpenStudy (anonymous):

initial amount

OpenStudy (anonymous):

yes so if A(t)= is the amount at time t, then A(0) will give you the initial amount of substance correct?

OpenStudy (anonymous):

so if you plug in t=0 to the equation then what do you get for an answer

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

\[A(0)=4(\frac{1}{2})^{\frac{1}{5}*0}=4(\frac{1}{2})^0\] what is anything to the zero power

OpenStudy (anonymous):

\[(\frac{1}{2})^0=?\]

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

yes so if that part is one then it's \[A(0)=4*1=4\]

OpenStudy (anonymous):

so it's saying that the amount at t =0 is 4, which just so happens to be what is in your function.

OpenStudy (anonymous):

so what does that mean that 4 is within the equation

OpenStudy (anonymous):

oops i'm sorry i thought they were the same question... but i gave you the answer for the first question A(t) is read as a "Function of time" , Amount of someting is a function therefore Amount of substance in terms of time

OpenStudy (anonymous):

or A(t) is the amount of substance at time t

OpenStudy (anonymous):

In that case would the second question be number of times to complete a cycle?

OpenStudy (anonymous):

for question two, is that suppose to be what does 1/6 mean

OpenStudy (anonymous):

yes i believe so

OpenStudy (anonymous):

Thanks